OpenStudy (anonymous):

Use continuity to evaluate the limit. lim x→π 5 sin(x + sin x)

OpenStudy (anonymous):

use the fact that sine is continuous what they want you to write is this \[\lim_{x\to \pi}5\sin(x+\sin(x))=5\lim_{x\to\pi}\sin(x+\sin(x))\] the first because the property that constants don't effect the limit then write \[=5\sin(\lim_{x\to \pi}(x+\sin(x))\] because sine is continuous, you can bring the limit inside the function more to come....

OpenStudy (anonymous):

next step \[=5\sin(\lim_{x\to \pi}x+\lim_{x\to \pi}\sin(x))\] because the limit of the sum is the sum of the limits

OpenStudy (anonymous):

next step \[=5\sin(\pi+\sin(\lim_{x\to \pi}x))\] because of two reasons, sine is continuous still so you can bring the limit inside, and also because \(\lim_{x\to \pi}x=\pi\) is obvious

OpenStudy (anonymous):

next step \[5\sin(\pi+\sin(\pi))\] for the same reason as before all this work is really just being silly, as everything is sight is continuous, you just replace all \(x\) you see by \(\pi\)

OpenStudy (anonymous):

final step is probably to evaluate since \(\sin(\pi)=0\) you get \[5\sin(\pi+0)=5\sin(\pi)=0\]

OpenStudy (anonymous):

thank you very much